OpenStudy (anonymous):

ln(x-6)=1 can anyone help?

OpenStudy (anonymous):

I suppose you are solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Suppose you have ln a = 2 You can take away the ln from a, but you must make 2 the exponent of e \[\ln a =2\]\[a =e ^{2}\]

OpenStudy (anonymous):

take e^ of both sides you get x-6= e^1 x= 6+e

OpenStudy (anonymous):

i think i understand, thank you.